OpenStudy (anonymous):

what is the maximums and minimums of the equation y=2.657x^3 + -1.48x^2 + .2819x + 7.7E-14?

OpenStudy (anonymous):

thats kinda complicated to figure out you should go to http://www.wolframalpha.com

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

if this is a y or f(x) then differentiate it then equaled it to zero and find its roots, that rot will be your max or min point..just plug it into your original equation....

OpenStudy (anonymous):

Y'=7.7971X^2 -2.96X +.2819.....NOW THIS IS A QUADRATIC EQUATION,.YOU CAN SOLVE FOR THE ROOTS NOW

OpenStudy (anonymous):

this is supposed to be a quartic...

OpenStudy (anonymous):

no,,, quartic eq starts at say x^4,....cubic is x^3.... quadratic is x^2

OpenStudy (anonymous):

what do i do after i derive the equation to find the minimums and maximums?

OpenStudy (anonymous):

this does start at x^4...... -1.28x4 + -33.36x3 + -323.41x2 + -1384.63x + -2206.60

OpenStudy (anonymous):

no it doesn't

OpenStudy (anonymous):

it's 2.657x^3.....

OpenStudy (anonymous):

hehe look up at your prob here atart at y=x^3 it a cubic eq rt?

OpenStudy (anonymous):

wait who are you talking to?

OpenStudy (anonymous):

after you done the derivative, the derivative is now a quadratic y'= ...x^2

OpenStudy (anonymous):

then how do you find the minimums and maximums

OpenStudy (anonymous):

ok ..you equate the derivative to zero and find its roots

OpenStudy (anonymous):

that roots will become you max or min of your original equation

OpenStudy (anonymous):

what is the answer after equating the derivative

OpenStudy (anonymous):

ok, to find the roots, do you know the quadratic formila?

OpenStudy (anonymous):

x=(-b+\[\sqrt{?}\]b2-4ac)2a

OpenStudy (anonymous):

im sorry my pc equation writer does not write well here you a=7.797.....b=-2.96...c= .2819

OpenStudy (anonymous):

x=(2.96+-\[\sqrt{?}\]-.0303)/15.594

OpenStudy (anonymous):

did you get it? the roots are imaginaries...do you know what is imaginary number?

OpenStudy (anonymous):

obviously... a number that's imaginary like i for example or e

OpenStudy (anonymous):

im just asking what are the points (max and min) for my equation

OpenStudy (anonymous):

no...any square root of a negative number are imaginary number like square root of -1, sqrot of -2, squroot of - numbers

OpenStudy (anonymous):

ok here the simplest way of geting the derivative roots is to graph the derivative and approximate its roots..did you get what i mean?

OpenStudy (anonymous):

Y'=7.7971X^2 -2.96X +.2819.....graph this and find its root or zeroes

OpenStudy (anonymous):

here one root is 0.2819...you plug it in your orig equation y=2.657(.2819)^3 + -1.48(.2819)^2 + .2819(.2819)+ 7.7E-14=?

OpenStudy (anonymous):

min y=0.02137744667

OpenStudy (anonymous):

i graph it rt now and found the root x=0.8948 now you can plug it in to your orig prob.

OpenStudy (anonymous):

your max or min y=0.9708296575

OpenStudy (anonymous):

sorry never mind that root x=.2819 or min y=.02137744667....thats a mistake,, graph it and found that the max or min y=.9708296575